OpenStudy (anonymous):

Need help will fan and medal let f(x)=-1/2x+5. the average rate of change from x=17 and x=43 equals? A.13 B.-1/2 C.1/2 D.-13

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

plug in 17 and 43 into x in the equation f(x). Ave rate of change = (f(43)-f(17)) / (43-17) which equates (-16.5 - (-3.5) ) / 26 = -13/26 = -1/2 or -0.5 Answer B

OpenStudy (anonymous):

 The rate of change from x=-10 to x=10 is ?